# Regret Minimization in Stackelberg Games with Side Information

### v3

@@ Abstract
Algorithms for playing in Stackelberg games have been deployed in real-world domains including airport security, anti-poaching efforts, and cyber-crime prevention. However, these algorithms often fail to take into consideration the additional information available to each player (e.g. traffic patterns, weather conditions, network congestion), a salient feature of reality which may significantly affect both players’ optimal strategies. We formalize such settings as Stackelberg games with side information, in which both players observe an external context before playing. The leader commits to a (context-dependent) strategy, and the follower best-responds to both the leader’s strategy and the context. We focus on the online setting in which a sequence of followers arrive over time, and the context may change from round-to-round. In sharp contrast to the non-contextual version, we show that it is impossible for the leader to achieve good performance (measured by regret) in the full adversarial setting. Motivated by our impossibility result, we show that no-regret learning is possible in two natural relaxations: the setting in which the sequence of followers is chosen stochastically and the sequence of contexts is adversarial, and the setting in which the sequence of contexts is stochastic and the sequence of followers is chosen by an adversary.
@@ 1 Introduction
Overview of our results.
In order to capture the additional information that the leader and follower may have at their disposal, we formalize such settings as Stackelberg games with side information. Specifically, we consider a setting in which a leader interacts with a sequence of followers in an online fashion. At each time-step, the leader gets to see payoff-relevant information about the current round in the form of a context. After observing the context, the leader commits to a mixed strategy over a finite set of actions, and the follower best-responds to both (1) the leader’s strategy and (2) the context in order to maximize their utility. We allow the follower in each round to be one of $K$ types. Each follower type corresponds to a different mapping from contexts, leader strategies, and follower actions to utilities. While the leader may observe the context before committing to their mixed strategy, they do not get to observe the follower’s type until after the round is over. Under this setting, the goal of the leader is to minimize their regret with respect to the best policy (i.e. the best mapping from contexts to mixed strategies) in hindsight, with respect to the realized sequence of followers and contexts.
@@
1.1 Related work
Learning in Stackelberg games.
Conitzer and Sandholm [9] provide algorithms for recovering the leader’s optimal mixed strategy in Stackelberg games when the follower’s payoff matrix is known to the leader. Letchford et al. [22] consider the problem of learning the leader’s optimal mixed strategy in the repeated Stackelberg game setting against a perfectly rational follower with an unknown payoff matrix. Peng et al. [24] study the same setting as Letchford et al. [22]. They provide improved rates and prove nearly-matching lower bounds. Learning algorithms to recover the leader’s optimal mixed strategy have also been studied in Stackelberg security games [5, 7, 27, 6].
Our work builds off of several results established for online learning in (non-contextual) Stackelberg games in Balcan et al. [5]. In particular, our results in Section 4.2 and Section C.2 may be viewed as a generalization of their results to the setting in which the payoffs of both players depend on an external context, in addition to the player actions. Roughly speaking, Balcan et al. [5] show that it without loss to play Hedge over a finite set of mixed strategies in order to obtain no-regret against an adversarially-chosen sequence of follower types. In order to handle the additional side information available in our setting, we instead play Hedge over a finite set of policies, each of which map to a finite set of (context-dependent) mixed strategies. However, the discretization argument is more nuanced in our setting. In particular, it is not without loss of generality to consider a finite set of policies. As a result, we need to bound the additional regret incurred by the leader due to the discretization. More recent work on learning in Stackelberg games provides improved regret rates in the full feedback [10] and bandit feedback [6] settings, and considers the effects of non-myopic followers [14] and followers who respond to calibrated predictions [15].
Side information in repeated game settings.
Lauffer et al. [21] study a Stackelberg game setting in which there is an underlying (probabilistic) state space which affects the leader’s rewards, and there is a single (unknown) follower type. In contrast, we study a setting in which the sequence of follower types and/or contexts may be chosen adversarially. As a result, their approach is generally not applicable in our setting. Sessa et al. [26] study a repeated game setting in which the players receive additional information (i.e. a context) at each round, much like in our setting. However, their focus is on repeated normal-form games, which require different tools and techniques to analyze compared to the repeated Stackelberg game setting we consider. Other work has also considered repeated normal-form games which change over time in different ways. In particular, Zhang et al. [34], Anagnostides et al. [3] study learning dynamics in time-varying game settings, and Harris et al. [17] study a meta-learning setting in which the game being played changes after a fixed number of rounds.
Contextual bandits.
@@ 2 Setting and background
Our setting.
We consider a game between a leader and a sequence of followers. In each round $t\in[T]$, Nature reveals a context $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{d}$ to both players. [^4] E.g. in airport security, $\mathbf{z}_{t}$ may contain information about arrival and departure times, number of passengers, valuable cargo, etc. In cyber-defense, $\mathbf{z}_{t}$ may be a list of network traffic statistics. The leader moves first by playing some mixed strategy $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{A}$ over a set of (finite) leader actions $\mathcal{A}$, i.e., $\mathbf{x}_{t}\in\Delta(\mathcal{A})$. The size of $\mathcal{A}$ is $A:=|\mathcal{A}|$. Having observed the leader’s mixed strategy, the follower best-responds to both $\mathbf{x}_{t}$ and $\mathbf{z}_{t}$ by playing some action $a_{f}\in\mathcal{A}_{f}$, where $\mathcal{A}_{f}$ is the (finite) set of follower actions and $A_{f}:=|\mathcal{A}_{f}|$.
@@
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the strategy given by the leader’s optimal-in-hindsight policy for context $\mathbf{z}$ is
$\pi^{*}(\mathbf{z})\in\arg\max_{\mathbf{x}\in\mathcal{X}}\sum_{t=1}^{T}u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))\cdot\mathbbm{1}\{\mathbf{z}_{t}=\mathbf{z}\}.$
We measure the leader’s performance against the optimal policy via the notion of contextual Stackelberg regret or (regret for short).
@@
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
where $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ is the sequence of mixed strategies played by the leader.
@@ 3 On the impossibility of fully adversarial no-regret learning
Online linear thresholding.
The online linear thresholding problem is a repeated two-player game between a learner and an adversary. Before the first round, an adversary chooses a cutoff $s\in[0,1]$ which is unknown to the learner. In each round, the adversary chooses a point $\omega_{t}\in[0,1]$ and reveals it to the learner. $\omega_{t}$ is assigned label $y_{t}=1$ if $\omega_{t}>s$ and label $y_{t}=-1$ otherwise. Given $\omega_{t}$, the learner makes a guess $g_{t}\in[0,1]$ (the probability they place on $y_{t}=1$), and receives utility $u_{\mathrm{OLT}}(\omega_{t},g_{t})=g_{t}\cdot\mathbbm{1}\{y_{t}=1\}+(1-g_{t})\cdot\mathbbm{1}\{y_{t}=-1\}$. The learner gets to observe $y_{t}$ after round $t$ is over. The learner’s policy $\pi_{t}:[0,1]\rightarrow[0,1]$ is a mapping from points in $[0,1]$ to guesses in $[0,1]$. The optimal policy $\pi^{*}$ makes guess $\pi^{*}(\omega_{t})=1$ if $\omega_{t}>s$ and $\pi^{*}(\omega_{t})=0$ otherwise. The learner’s regret after $T$ rounds is given by $R_{\mathrm{OLT}}(T)=T-\sum_{t=1}^{T}u_{\mathrm{OLT}}(\omega_{t},g_{t})$, since the optimal policy achieves utility $1$ in every round. In order to prove a lower bound on contextual Stackelberg regret in our setting, we make use of the following well-known lower bound on regret in the online linear thresholding setting (see e.g. [13]).
@@ 4.1 Stochastic follower types and adversarial contexts
Given a distribution over followers $\mathcal{F}$ and a sequence of contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the leader’s expected contextual Stackeleberg regret is
$\mathbb{E}[R(T)]:=\mathbb{E}_{f_{1},\ldots,f_{T}\sim\mathcal{F}}\left[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right],$
where $\pi^{*}$ is the optimal policy given knowledge of $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and $\mathcal{F}$.
Under this setting, the utility for policy $\pi$ may be written as
$\mathbb{E}_{f_{1},\ldots,f_{T}\sim\mathcal{F}}\left[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))\right]=\sum_{t=1}^{T}\mathbb{E}_{f_{1},\ldots,f_{t-1}\sim\mathcal{F}}[\mathbb{E}_{t}[u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))]].$
Our algorithm (Algorithm 1) proceeds by estimating the inner expectation $\mathbb{E}_{t}[u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))]$ as
@@
Input: $\widehat{\mathbf{p}}_{t}$

### v4

@@ Abstract
Algorithms for playing in Stackelberg games have been deployed in real-world domains including airport security, anti-poaching efforts, and cyber-crime prevention. However, these algorithms often fail to take into consideration the additional information available to each player (e.g. traffic patterns, weather conditions, network congestion), which may significantly affect both players’ optimal strategies. We formalize such settings as Stackelberg games with side information, in which both players observe an external context before playing. The leader commits to a (context-dependent) strategy, and the follower best-responds to both the leader’s strategy and the context. We focus on the online setting in which a sequence of followers arrive over time, and the context may change from round-to-round. In sharp contrast to the non-contextual version, we show that it is impossible for the leader to achieve no-regret in the full adversarial setting. Motivated by this result, we show that no-regret learning is possible in two natural relaxations: the setting in which the sequence of followers is chosen stochastically and the sequence of contexts is adversarial, and the setting in which contexts are stochastic and follower types are adversarial.
@@ 1 Introduction
Overview of our results. In order to capture the additional information that the leader and follower may have at their disposal, we formalize such settings as Stackelberg games with side information. Specifically, we consider a setting in which a leader interacts with a sequence of followers in an online fashion. At each time-step, the leader gets to see payoff-relevant information about the current round in the form of a context. After observing the context, the leader commits to a mixed strategy over a finite set of actions, and the follower best-responds to both (1) the leader’s strategy and (2) the context in order to maximize their utility. We allow the follower in each round to be one of $K$ types. Each follower type corresponds to a different mapping from contexts, leader strategies, and follower actions to utilities. While the leader may observe the context before committing to their mixed strategy, they do not get to observe the follower’s type until after the round is over. Under this setting, the goal of the leader is to minimize their regret with respect to the best policy (i.e. the best mapping from contexts to mixed strategies) in hindsight, with respect to the realized sequence of followers and contexts.
@@
Related work. Letchford et al. [21] consider the problem of learning the leader’s optimal mixed strategy in the repeated Stackelberg game setting against a perfectly rational follower with an unknown payoff matrix. Peng et al. [23] study the same setting as Letchford et al. [21]. They provide improved rates and prove nearly-matching lower bounds. Learning algorithms to recover the leader’s optimal mixed strategy have also been studied in Stackelberg security games [5, 7, 26, 6].
Our work builds off of several results established for online learning in (non-contextual) Stackelberg games in Balcan et al. [5]. In particular, our results in Section 4.2 and Section C.2 may be viewed as a generalization of their results to the setting in which the payoffs of both players depend on an external context. Roughly speaking, Balcan et al. [5] show that it without loss to play Hedge over a finite set of mixed strategies in order to obtain no-regret against an adversarially-chosen sequence of follower types. In order to handle the additional side information available in our setting, we instead play Hedge over a finite set of policies, each of which map to a finite set of (context-dependent) mixed strategies. However, the discretization argument is more nuanced in our setting. In particular, it is not without loss of generality to consider a finite set of policies. As a result, we need to bound the additional regret incurred by the leader due to the discretization. More recent work on learning in Stackelberg games provides improved regret rates in the full feedback [9] and bandit feedback [6] settings, and considers the effects of non-myopic followers [13] and followers who respond to calibrated predictions [14].
Lauffer et al. [20] study a Stackelberg game setting in which there is an underlying (probabilistic) state space which affects the leader’s rewards, and there is a single (unknown) follower type. In contrast, we study a setting in which the sequence of follower types and/or contexts may be chosen adversarially. Sessa et al. [25] study a repeated game setting in which the players receive additional information (i.e. a context) at each round, much like in our setting. However, their focus is on repeated normal-form games, which require different tools and techniques to analyze compared to the repeated Stackelberg game setting we consider. Other work has also considered repeated normal-form games which change over time in different ways. In particular, Zhang et al. [33], Anagnostides et al. [3] study learning dynamics in time-varying game settings, and Harris et al. [16] study a meta-learning setting in which the game being played changes after a fixed number of rounds.
@@ 2 Setting and background
Our setting. We consider a game between a leader and a sequence of followers. In each round $t\in[T]$, Nature reveals a context $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{d}$ to both players. [^4] E.g. in airport security, $\mathbf{z}_{t}$ may contain information about arrival and departure times, number of passengers, valuable cargo, etc. In cyber-defense, $\mathbf{z}_{t}$ may be a list of network traffic statistics. The leader moves first by playing some mixed strategy $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{A}$ over a set of (finite) leader actions $\mathcal{A}$, i.e., $\mathbf{x}_{t}\in\Delta(\mathcal{A})$. The size of $\mathcal{A}$ is $A:=|\mathcal{A}|$. Having observed the leader’s mixed strategy, the follower best-responds to both $\mathbf{x}_{t}$ and $\mathbf{z}_{t}$ by playing some action $a_{f}\in\mathcal{A}_{f}$, where $\mathcal{A}_{f}$ is the (finite) set of follower actions and $A_{f}:=|\mathcal{A}_{f}|$.
@@
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the strategy given by the leader’s optimal-in-hindsight policy for context $\mathbf{z}$ is $\pi^{*}(\mathbf{z})\in\arg\max_{\mathbf{x}\in\mathcal{X}}\sum_{t=1}^{T}u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))\cdot\mathbbm{1}\{\mathbf{z}_{t}=\mathbf{z}\}$.
We measure the leader’s performance against the optimal policy via the notion of contextual Stackelberg regret (regret for short).
@@
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is $R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ is the sequence of mixed strategies played by the leader.
@@ 3 On the impossibility of fully adversarial no-regret learning
Online linear thresholding. The online linear thresholding problem is a repeated two-player game between a learner and an adversary. Before the first round, an adversary chooses a cutoff $s\in[0,1]$ which is unknown to the learner. In each round, the adversary chooses a point $\omega_{t}\in[0,1]$ and reveals it to the learner. $\omega_{t}$ is assigned label $y_{t}=1$ if $\omega_{t}>s$ and label $y_{t}=-1$ otherwise. Given $\omega_{t}$, the learner makes a guess $g_{t}\in[0,1]$ (the probability they place on $y_{t}=1$), and receives utility $u_{\mathrm{OLT}}(\omega_{t},g_{t})=g_{t}\cdot\mathbbm{1}\{y_{t}=1\}+(1-g_{t})\cdot\mathbbm{1}\{y_{t}=-1\}$. The learner gets to observe $y_{t}$ after round $t$ is over. The learner’s policy $\pi_{t}:[0,1]\rightarrow[0,1]$ is a mapping from points in $[0,1]$ to guesses in $[0,1]$. The optimal policy $\pi^{*}$ makes guess $\pi^{*}(\omega_{t})=1$ if $\omega_{t}>s$ and $\pi^{*}(\omega_{t})=0$ otherwise. The learner’s regret after $T$ rounds is given by $R_{\mathrm{OLT}}(T)=T-\sum_{t=1}^{T}u_{\mathrm{OLT}}(\omega_{t},g_{t})$, since the optimal policy achieves utility $1$ in every round. In order to prove a lower bound on contextual Stackelberg regret in our setting, we make use of the following well-known lower bound on regret in the online linear thresholding setting (see e.g. [12]).
@@ 4.1 Stochastic follower types and adversarial contexts
Given a distribution over followers $\mathcal{F}$ and a sequence of contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$, the leader’s expected contextual Stackeleberg regret is $\mathbb{E}[R(T)]:=\mathbb{E}_{f_{1},\ldots,f_{T}\sim\mathcal{F}}[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))]$, where $\pi^{*}$ is the optimal policy given knowledge of $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and $\mathcal{F}$.
Under this setting, the utility for policy $\pi$ may be written as $\mathbb{E}_{f_{1},\ldots,f_{T}\sim\mathcal{F}}\left[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))\right]=\sum_{t=1}^{T}\mathbb{E}_{f_{1},\ldots,f_{t-1}\sim\mathcal{F}}[\mathbb{E}_{t}[u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))]]$. Our algorithm (Algorithm 1) proceeds by estimating the inner expectation $\mathbb{E}_{t}[u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))]$ as
@@
Input: $\widehat{\mathbf{p}}_{1}$
@@ 4.2 Stochastic contexts and adversarial follower types
4.3 Simulations
(a) Non-stochastic contexts, stochastic follower types.
(b) Stochastic contexts, non-stochastic follower types.
(c) Stochastic contexts, stochastic follower types.
Figure 2: Cumulative average reward of Algorithm 1, Algorithm 2, and the algorithm of Balcan et al. [5] (which does not take side information into consideration) over five runs in a synthetic data setup. Shaded regions represent one standard deviation.
We empirically evaluate the performance of Algorithm 1 and Algorithm 2 on synthetically-generated data. We consider a setup in which $K=5$, $A=A_{f}=3$, and the context dimension $d=3$. Utility functions are linear in both the context and player actions, and are sampled u.a.r. from $[-1,1]^{3\times 3\times 3}$.
We compare the cumulative reward of our algorithms to each other and the algorithm of Balcan et al. [5] (which does not leverage side information) as a baseline. We simulate non-stochastic context arrivals in Figure 2(a) by displaying the same context for $T/4$ time-steps in a row. Follower types are chosen u.a.r. from each of the five follower types. In Figure 2(b), contexts are generated stochastically by sampling each component u.a.r. from $[-1,1]$. Followers are chosen non-stochastically by deterministically cycling over the five types. In Figure 2(c), both contexts and follower types are chosen stochastically. Specifically, contexts are generated as in Figure 2(b) and follower types are generated as in Figure 2(a).
We find that Algorithm 1 and Algorithm 2 perform similarly across instances, and both significantly out-perform the baseline of Balcan et al. [5]. It would be interesting to find instances for which Algorithm 1 (resp. Algorithm 2) performs poorly whenever followers (resp. contexts) are chosen non-stochastically.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by NSF Grants CCF-1910321, #1763786, and by an NDSEG fellowship. The authors would like to thank the anonymous reviewers for valuable feedback, and Martino Bernasconi, Matteo Castiglioni, and Andrea Celli for helpful discussions surrounding related work.